OpenStudy (anonymous):

solve for x.x on the outside of sqrt of 2+7=sqrt 3

OpenStudy (anonymous):

As in: \(x\sqrt{2+7}=\sqrt{3}\)?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

\[x \sqrt{2+7}=x \sqrt{9}=\sqrt{3}\implies x=\frac{\sqrt{3}}{\sqrt{9}}=\sqrt{\frac{3}{9}}=\sqrt{\frac{1}{3}}=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

that is wrong your possible answers are A.sqrt of 6-7 sqrt of 2 over 2 B:sqrt of 6-7 sqrt of 2 over 4 C.sqrt 3-7 sqrt of 2 over 2

OpenStudy (anonymous):

In that case you told me the question wrong. It's A.

OpenStudy (anonymous):

no i didnt i wrote it exactly how it said

OpenStudy (anonymous):

I asked you if it was: \(x\sqrt{2+7}=\sqrt{3}\) and you said yes. It's actually: \(x\sqrt{2}+7=\sqrt{3}\).

OpenStudy (anonymous):

ohhhhhhhhhhhh

OpenStudy (anonymous):

how do you make the sqrt sign?

OpenStudy (anonymous):

There's a button for it in equation editor on the 4th tab down but if you want raw latex, it's '\sqrt'.